# Angles Don’t Lie: Unlocking Training-Efficient RL Through the Model’s Own Signals

### v1

@@ B.2 Attention-Based Explanation of Layer-wise Angle Concentration Patterns
$\begin{split}\alpha_{im}&=Softmax\bigl{(}(\hat{y}_{i}\,W_{q})\,\bigl{(}\hat{y}_{j}\,W_{k}\bigr{)}^{T}/\sqrt{d}\bigr{)}=Softmax\bigl{(}\langle\hat{y}_{i}W_{q},\hat{y}_{j}W_{k}\rangle/\sqrt{d}\bigr{)}\\ &=Softmax\bigl{(}\hat{y}_{i}(W_{q}W_{k}^{T})\hat{y}_{j}^{T}/\sqrt{d}\bigr{)}=Softmax\bigl{(}\theta\cdot\langle\hat{y}_{i},\hat{y}_{j}\rangle/\sqrt{d}\bigr{)}\\ \end{split}$ (18)
@@
$\begin{split}\alpha_{im}=Softmax\bigl{(}\theta\cdot\cos(\angle(y_{i},y_{j})/\sqrt{d}\bigr{)}\propto\cos(\angle(y_{i},y_{j})\bigr{)}\end{split}$ (19)
@@ C.2 Data-wise Angle Concentration Patterns
Fig. 19 provides a more fine-grained view of the data-wise angle-concentration patterns. Consistent with the conclusions in the main text, it reveals a curriculum-like trend: the model learns samples exhibiting high angular concentration earlier, followed by samples with lower angular concentration.
@@ 16             Record $\mathcal{C}_{M_{t}}(d^{(t)}_{i})$ and accuracy $\mathrm{Acc}_{M_{t}}(d^{(t)}_{i})$
18      Update $M_{t}$ with accuracy-based loss
@@ 20       Compute mean angle $\mathcal{C}^{(t)}=\frac{1}{n}\sum_{i=1}^{n}\mathcal{C}_{M_{t}}(d^{(t)}_{i})$
21       Update sampling-mean
@@
22
Algorithm 1 GAIN–RL (GRPO): Gradient-driven Angle-Informed Navigated RL Framework
@@ F.2 Model Performance on Single Task
Fig. 9 in the main text illustrates the training dynamics of Qwen-2.5-0.5B-Instruct on three single-task datasets. For a more detailed comparison, Tab. LABEL:tab_single reports the final performance at epoch 200 and the corresponding training speedup across different training sets and methods. Notably, on the GSM8K dataset, GAIN-RL outperforms the original GRPO by 4.72% in final accuracy and achieves a 3.3$\times$ improvement in training speed.
@@
Table 5: Fine-tuning performance on single tasks. Models are trained on the training sets and evaluated on their validation sets. ADARFT is excluded on GSM8K due to missing difficulty coefficients.
@@
GRPO - - 48.43 200 1$\times$ 34.80 200 1$\times$ 9.64 200 1$\times$
@@
GAIN-RL(GRPO) Angle $<$ 10 min 53.15 60 3.33$\times$ 37.40 80 2.50$\times$ 12.04 100 2.00$\times$

### v2

@@ B.2 Attention-Based Explanation of Layer-wise Angle Concentration Patterns
[…]
@@
$\begin{split}\alpha_{ij}=Softmax\bigl(\theta\cdot\cos(\angle(y_{i},y_{j})/\sqrt{d}\bigr)\propto\cos(\angle(y_{i},y_{j})\bigr)\end{split}$ (19)
@@ C.2 Data-wise Angle Concentration Patterns
Fig. 19 provides a more fine‑grained view of the data‑wise angle‑concentration patterns. Consistent with the conclusions in the main text, it reveals a curriculum‑like trend: the model learns samples exhibiting high angular concentration earlier, followed by samples with lower angular concentration.
@@ 16      Record $\mathcal{C}_{M_{t}}(d^{(t)}_{i})$ and accuracy $\mathrm{Acc}_{M_{t}}(d^{(t)}_{i})$
18  Update $M_{t}$ with accuracy‑based loss
@@ 20   Compute mean angle $\mathcal{C}^{(t)}=\frac{1}{n}\sum_{i=1}^{n}\mathcal{C}_{M_{t}}(d^{(t)}_{i})$
21   Update sampling‐mean
22  -0.35em
@@ 21   Update sampling‐mean
23
Algorithm 1 GAIN–RL (GRPO): Gradient-driven Angle‑Informed Navigated RL Framework
@@ F.2 Model Performance on Single Task
[…]
@@
Table 5: Fine‑tuning performance on single tasks. Models are trained on the training sets and evaluated on their validation sets. ADARFT is excluded on GSM8K due to missing difficulty coefficients.
@@
[2pt/2pt] ADARFT(GRPO) Difficulty
>> 1 day
@@
[…]